Question 8 Try again. Hint: More than four (4) means (1- Probability (up to and including four (4)). Traffic engineers in
Answer #1

a) here pth percentile =-ln(1-p)/\lambda=-ln(1-0.15)/\lambda =50040

\lambda =0.00000325 faiilure/hour

a)

expected number of failure in 50040 hours=50040*0.00000325 =0.162519

hence P(fewer than 2 bulb fail) =P(X<=1)=P(X=0)+P(X=1)=0.988

b)P(no bulb fail)=P(X=0)=e-0.162519*0.1625190/0! =0.85

c)P(X>4)=1-P(X<=4)=1-0.999999 =0.000001